Your firm is considering a project that will cost $4.548 million up​ front, generate cash flows of $3.50 million per year for 3 ​years, and then have a cleanup and shutdown cost of $6.00 million in the fourth year.

a. How many IRRs does this project​ have?

b. Calculate a modified IRR for this project assuming a discount and compounding rate of 10.0%.

c. Using the MIRR and a cost of capital of 10.0%​,

would you take the​ project?

a. How many IRRs does this project​ have?

The project has

1

2

3

4

IRRs.  ​(Select from the​ drop-down menu.)
Answer #1

a. There are 2 irrs because there is change of sign cash flows twice.

b. PV of cost =4.548+6/1.1^4 =8.6461
FV of cash flows =3.50*1.1^3+3.50*1.1^2+3.50*1.1 =12.7435
MIRR=(12.7435/8.6461)^(1/4)-1=10.18%

c. Since MIRR is greater than cost of capital project should be accepted